OpenStudy (anonymous):

If my bro is 3 years older than me and my sis is 4 years younger than me and the product of the ages is 1872 then how old is my sister? Something about polynomial equations If my bro is 3 years older than me and my sis is 4 years younger than me and the product of the ages is 1872 then how old is my sister? Something about polynomial equations @Mathematics

OpenStudy (anonymous):

(x+3)*(x-4) = 1872

OpenStudy (anonymous):

oh...okay

OpenStudy (anonymous):

x is ur age ur sister age is (x-4)

OpenStudy (anonymous):

right but that puts me at x= 1876 :/

OpenStudy (mathteacher1729):

I believe the equation should be written as: x(x-4)(x+3) = 1872 x = your age (x-4) = "Sister is four year younger than you" (x+3) = "Bro is three years younger than you" "the product of the ages is 1872" means "multiply all the ages together" Also, the equation x(x-4)(x+3) = 1872 has THREE roots, which will give three ages. Also the age x is a real number that makes sense. :D

OpenStudy (mathteacher1729):

(x+3) = "Bro is three years OLDER than you" * typo sorry

OpenStudy (anonymous):

Yay thank you!

OpenStudy (mathteacher1729):

Point of clarification: "Also, the equation x(x-4)(x+3) = 1872 has THREE roots, which will give three ages. Also the age x is a real number that makes sense. :D" What I should have said here is that the equation has one REAL solution (which makes sense in the context of this problem). Once you obtain that real solution, you can plug it back into the equation to get Sis and Bro's age. The other two solution obviously don't matter, since they are COMPLEX (they involve sqrt(-1) = i, which doesn't make sense, cuz ages are real numbers).